OpenStudy (anonymous):

Which of the following are solutions to the equation below? Check all that apply. x2 - 2x = 24

OpenStudy (anonymous):

A. x = 6 B. x = -4 C. x = -3 D. x = -6 E. x = 4 F. x = 3

OpenStudy (anonymous):

how do i do this problem?

OpenStudy (anonymous):

do i just plug x in?

OpenStudy (hoblos):

yes you can just plug x in and see if the equation is true plug in 6 6^2 - 2(6) = 36 - 12 = 24 true plug in -4 (-4)^2 - 2(-4) = 16+8 = 24 true you can also try to factorize the equation if you know it would be x2 - 2x = 24 x2 - 2x -24 = 0 (x-6)(x+4)= 0 so x=6 or x=-4

OpenStudy (anonymous):

Alright, thanks.

OpenStudy (hoblos):

glad to help :)